OpenStudy (anonymous):

Parametric equations. Find equation of tangent line to the curve x=e^t and y=e^(-t) by eliminating the parameter

OpenStudy (anonymous):

There should be a point (a,b) given to find an equation.

OpenStudy (anonymous):

Start by writing everything in terms of t; x=e^t ==> t = ln(x) and y = e^-(lnx) and then take the derivative of y with respect to x, that should be the answer